# SplashNet: Split-and-Share Encoders for Accurate and Efficient Typing with Surface Electromyography

### v1

@@ 5.2 Finetuned Model Performance
We next evaluate the performance of models finetuned on user-specific data. As in Sivakumar et al. (14), we maintain identical training hyperparameters during both generic pretraining and finetuning. While this simplifies the analysis, we note that the optimal hyperparameters for finetuning likely differ from those used during pretraining, and further gains may be achievable through phase-specific hyperparameter tuning.
@@
[…]
We next examine architectural changes that explicitly encode inductive biases about the bilateral and causal structure of EMG typing. Both the Split-only and SplashNet-mini models yield substantial improvements over the Joint-Hand baseline, with CERs of 6.10% and 5.87% ($p=6e$-$3$ and $p=2e$-$3$, respectively). For the latter, we evaluate two strategies during finetuning: either maintaining shared weights or duplicating them to allow separate adaptation per hand. Interestingly, both strategies yield similar performance, suggesting that hand-specific encoder weights are unnecessary even in the finetuned setting.
[…]
@@
Method Test domain val Test domain test GFLOPs (30 s) Params
Sivakumar et al. 2024 8.31 $\pm$ 3.19 6.95 $\pm$ 3.61 201.325 5.29M
+ RSG 6.70 $\pm$ 3.22 6.92 $\pm$ 3.79 178.948 4.96M
+ RTN 6.47 $\pm$ 2.92 6.63 $\pm$ 3.58 178.948 4.96M
+ ACM 7.18 $\pm$ 3.33 7.45 $\pm$ 3.87 178.948 4.96M
+ RTN + ACM 6.55 $\pm$ 2.91 6.53 $\pm$ 3.27 178.948 4.96M
+ Split 6.04 $\pm$ 2.81 6.10 $\pm$ 3.32 118.350 2.68M
      + Share (SplashNet-mini) 6.13 $\pm$ 2.96 5.87 $\pm$ 3.04 118.350 1.38M
       + FT Unshare 5.85 $\pm$ 2.83 5.96 $\pm$ 3.28 118.350 2.68M
       + Upscale (SplashNet) 5.46 $\pm$ 2.60 5.51 $\pm$ 2.81 229.384 2.58M
        + FT Unshare 5.57 $\pm$ 2.65 5.67 $\pm$ 2.97 229.384 5.06M
@@ 6 Discussion
The central contribution of this work is instilling simple, well-motivated priors through preprocessing, augmentation and architecture that can close a surprising fraction of the generalization gap in wrist-EMG typing. SplashNet-mini and SplashNet both achieve large absolute and relative CER reductions (-$\mbox{-}15.4$ pp / -$\mbox{-}31.1\%$ zero-shot; -$\mbox{-}1.44$ pp / -$\mbox{-}20.7\%$ after fine-tuning) while cutting parameters to ¼–½ of the baseline and FLOPs to 0.6–1.15×. These gains are on par with (and complementary to) the $\sim 25\%$ error reduction CTRL-Labs at Reality Labs et al. (2) reported from doubling dataset size in a handwriting task, suggesting that principled inductive biases are as potent as raw data scaling for sEMG.
@@
A practical ambition articulated in Sivakumar et al. (14) is to run the entire decoder on the wristbands themselves, thereby mitigating concerns around latency, privacy, and robustness to Bluetooth interference. Achieving on-band inference will require an architecture that is not only light-weight but also fully split—able to process and output keystrokes from each wrist independently, without any cross-hand coupling. SplashNet moves part-way toward this goal by duplicating (and sharing) the encoder streams, yet it still merges information at the final linear output layer, so each wrist must communicate its embeddings to the other. Removing this last dependency, or replacing it with a low-bandwidth handshake, remains an open engineering challenge and a fruitful direction for future model architecture work.
@@
Beyond these concrete results, our study points to several avenues. Normalization: RTN is a single-pass causal z-score; richer adaptive schemes (momentum updates, learnable affine transforms, or brief self-supervised calibration) may yield further robustness. Structured masking: ACM already improves zero-shot CER, but spatially contiguous electrode “drop-blocks”, global time masks, or anatomical adjacency priors could guide the network toward even more transferable features. Model scale and self-supervision: SplashNet shows that capacity can be reinvested profitably once good priors are in place; coupling larger split-and-share encoders with masked-prediction pre-training may unlock still-higher accuracy without additional labels.
@@
Two limitations deserve emphasis. First, our models are developed for the practical beam-search + LM pipeline; under pure greedy decoding, the Joint-Hand baseline scores the best on unseen users, and ACM dents fine-tuned accuracy (details and discussion in Appendix A.3). Second, all experiments use healthy participants; the extent to which our priors transfer to populations with motor or limb differences (where electrode placement and muscle recruitment differ markedly) remains to be tested.
@@
In sum, we show that the long-standing vision of “keyboard-quality” EMG typing can be advanced not only by more data but also by better assumptions. By pairing causal normalization, aggressive yet structured regularization, and a symmetry-aware encoder, we make significant strides towards truly out-of-the-box wrist-sEMG typing—paving the way for real-world assistive and AR/VR interfaces.
@@ A.2 Evaluation Details
A.3 Greedy vs. Beam-Search Decoding
Our main text focuses on beam-search decoding with an external 6-gram language model (LM), because this is the configuration most relevant to realistic, latency-constrained deployment. For completeness, we also report greedy CTC decoding results,i.e. decoding without any LM.
Table 3: Zero-shot CER (%) under greedy and beam-search decoding.
Method Other domain val (greedy) Other domain val (beam) Test domain val (greedy) Test domain val (beam) Test domain test (greedy) Test domain test (beam)
Sivakumar et al. 2024 55.57 52.10 55.38 51.78
+ RSG 68.55 67.48 52.24 47.26 52.27 47.18
     + RTN 64.19 61.95 46.31 39.49 46.07 39.15
     + ACM 65.60 63.08 48.87 42.62 49.23 42.62
     + RTN + ACM 61.82 58.85 43.72 36.41 43.80 36.42
       + Split 61.74 58.64 45.73 37.28 45.69 37.37
        + Share (SplashNet-mini) 61.07 58.20 45.33 36.46 45.26 36.41
          + Upscale (SplashNet) 60.16 56.95 44.79 35.49 44.78 35.67
Zero-shot generalization.
Table 3 shows that, in the zero-shot regime, the joint-hand architecture with RTN and ACM often surpasses the Split-only and Split-and-Share architectures under greedy decoding, even though the ranking reverses once beam search is applied. The most plausible explanation is that the joint-hand encoder can observe both wrists simultaneously and therefore learns a stronger implicit LM, capturing bi- and tri-gram dependencies spread across hands. While that emergent linguistic prior is helpful when no external LM is available, it becomes redundant—and potentially counter-productive—once a more reliable 6-gram LM is introduced at inference time.
Table 4: Fine-tuned CER (%) with and without beam search.
Method Test domain val (greedy) Test domain val (beam) Test domain test (greedy) Test domain test (beam)
@@
Fine-tuned models.
A different pattern emerges after user-specific fine-tuning (Table 4). None of our variants fully match the baseline of Sivakumar et al. [2024] under greedy decoding, despite decisive gains once beam search is enabled. Because the only changes between the baseline and the RSG + 16 s variant are (i) four-fold longer clips during training, (ii) coarser spectral resolution, and (iii) a five-fold higher probability of channel masking, the degradation must stem from one (or a combination) of these factors.
Two clues implicate aggressive channel masking (ACM). First, the model with RTN but not ACM except for ACM achieves the best greedy scores, whereas the model with just ACM and not RTN yields the worst greedy performance. Second, the performance gap between the model with just RTN and the model with both RTN and ACM vanishes under beam search, indicating that the external LM compensates for information lost when ACM forces the network to rely on very low-order feature combinations. We therefore hypothesize that ACM, while beneficial for cross-user generalization, removes within-user cues that help distinguish confusable keystrokes, a weakness that beam-search decoding can largely recover.
In summary, greedy decoding accentuates two complementary inductive biases: (1) joint-hand encoders learn a useful internal LM, an advantage that vanishes once an external LM is applied, and (2) aggressive masking of input channels trades cross-user generalization for within-user discriminability, a trade-off that RTN and beam-search decoding can effectively offset.
@@ A.4 Calculation of FLOPs
A.5 Backspace-aware beam search with 6-gram Character LM
For decoding we adopt the backspace-aware beam search of Sivakumar et al. [2024], which keeps the 50 most-probable CTC label prefixes at each frame. Blank transitions simply update the score of an unchanged prefix, while the highest-scoring non-blank characters extend each hypothesis and are re-ranked with a 6-gram character LM prior. When a back-space symbol appears, the algorithm retracts the last language-model contribution, letting deletions correct earlier mistakes without extra passes. After the final frame any open LM context is closed, and the best-scoring prefix provides the decoded keystroke sequence. All beam-search parameters are the same as in Sivakumar et al. [2024]
A.6 Compute Resources

### v2

@@ 5.2 Finetuned Model Performance
Method
Test
domain
val
Test
domain
test
GFLOPs
(30 s)
Params
Sivakumar et al. 2024 8.31 $\pm$ 3.19 6.95 $\pm$ 3.61 61.61 5.29M
+ RSG 6.70 $\pm$ 3.22 6.92 $\pm$ 3.79 54.15 4.96M
+ RTN 6.47 $\pm$ 2.92 6.63 $\pm$ 3.58 54.15 4.96M
+ ACM 7.18 $\pm$ 3.33 7.45 $\pm$ 3.87 54.15 4.96M
+ RTN + ACM 6.55 $\pm$ 2.91 6.53 $\pm$ 3.27 54.15 4.96M
+ Split 6.04 $\pm$ 2.81 6.10 $\pm$ 3.32 36.84 2.68M
    + Share (SplashNet-mini) 6.13 $\pm$ 2.96 5.87 $\pm$ 3.04 36.84 1.38M
    + FT Unshare 5.85 $\pm$ 2.83 5.96 $\pm$ 3.28 36.84 2.68M
    + Upscale (SplashNet) 5.46 $\pm$ 2.60 5.51 $\pm$ 2.81 71.38 2.58M
     + FT Unshare 5.57 $\pm$ 2.65 5.67 $\pm$ 2.97 71.38 5.06M
We next evaluate the performance of models finetuned on user-specific data. As in Sivakumar et al. (14), we maintain identical training hyperparameters during both generic pretraining and finetuning. While this simplifies the analysis, we note that the optimal hyperparameters for finetuning likely differ from those used during pretraining, and further gains may be achievable through phase-specific hyperparameter tuning. We also note that, because each recording session involved doffing and re-donning the wristbands, the finetuning experiments inherently probe generalization across electrode placements in different sessions.
@@
First, we assess whether the same methods that improved zero-shot generalization also enhance performance in the finetuned setting. Using RSG alone yields similar performance to the baseline of Sivakumar et al. (14), with a CER of 6.91%, while adding RTN yields a significantly improved CER of 6.63% ($p=.028$ vs. +RSG). Interestingly, applying ACM in isolation—without RTN—leads to worse performance than the baseline of Sivakumar et al. (14), whereas the Joint-Hand baseline model, with both ACM and RTN, matches the model with RTN alone (6.53% CER). This suggests that masking-induced variability may hinder learning when the model lacks an appropriate normalization strategy to stabilize the input feature space. Moreover, unlike the generic case, finetuned models may rely more on higher-order feature correlations that are relatively stable across sessions for the same user, diminishing the benefits of aggressive masking.
We next examine architectural changes that explicitly encode inductive biases about the bilateral and causal structure of EMG typing. Both the Split-only and SplashNet-mini models yield substantial improvements over the Joint-Hand baseline, with CERs of 6.10% and 5.87% (p=6ep=6e-33 and p=2ep=2e-33, respectively). For the latter, we evaluate two strategies during finetuning: either maintaining shared weights or duplicating them to allow separate adaptation per hand. Interestingly, both strategies yield similar performance, suggesting that hand-specific encoder weights are unnecessary even in the finetuned setting, although it is possible that unsharing weights could become advantageous with more user-specfic data for finetuning.
Finally, we evaluate the SplashNet model. Again, we do not find any benefit from unsharing the weights during finetuning. With the weights kept shared, SplashNet achieves the best performance overall with a CER of 5.51% ($p=.02$ vs. SplashNet-mini), establishing a new state-of-the-art for user-specific EMG keystroke decoding in this benchmark.
@@ 6 Discussion
The central contribution of this work is instilling simple, well-motivated priors through preprocessing, augmentation and architecture that can close a surprising fraction of the generalization gap in wrist‑EMG typing. SplashNet‑mini and SplashNet both achieve large absolute and relative CER reductions (15.4‑15.4 pp / $31.1\%$ zero‑shot; 1.44‑1.44 pp / $20.7\%$ after fine‑tuning) while cutting parameters to ¼–½ of the baseline and FLOPs to 0.6–1.15×. These gains are on par with (and complementary to) the $\sim 25\%$ error reduction CTRL-Labs at Reality Labs et al. (2) reported from doubling dataset size in a handwriting task, suggesting that principled inductive biases are as potent as raw data scaling for sEMG.
@@
A practical ambition articulated in Sivakumar et al. (14) is to run the entire decoder on the wristbands themselves, thereby mitigating concerns around latency, privacy, and robustness to Bluetooth interference. Achieving on‑band inference will require an architecture that is not only light‑weight but also fully split—able to process and output keystrokes from each wrist independently, without any cross‑hand coupling. SplashNet moves part‑way toward this goal by duplicating (and sharing) the encoder streams, yet it still merges information at the final linear output layer, so each wrist must communicate its embeddings to the other. Removing this last dependency, or replacing it with a low‑bandwidth handshake, remains an open engineering challenge and a fruitful direction for future model architecture work. Another promising direction is exploring hybrid cross-hand architectures that retain split-and-share components but introduce limited, structured interactions between the hands, potentially offering a middle ground between the Split-and-Share and Joint-Hand extremes.
@@
Beyond these concrete results, our study points to several avenues. Normalization: RTN is a single‑pass causal z‑score; richer adaptive schemes (momentum updates, learnable affine transforms, or brief self‑supervised calibration) may yield further robustness. Structured masking: ACM already improves zero‑shot CER, but spatially contiguous electrode “drop‑blocks”, global time masks, or anatomical adjacency priors could guide the network toward even more transferable features. Model scale and self‑supervision: SplashNet shows that capacity can be reinvested profitably once good priors are in place; coupling larger split‑and‑share encoders with masked‑prediction pre‑training may unlock still‑higher accuracy without additional labels.
@@
Two limitations deserve emphasis. First, our models are developed for the practical beam-search + LM pipeline; under pure greedy decoding, the Joint-Hand baseline scores the best on unseen users, and ACM dents finetuned accuracy (details and discussion in Appendix A.3). This likely reflects the fact that the Joint-Hand baseline can exploit cross-hand co-articulatory patterns, which tend to correlate strongly with character bi- and tri-grams. These weak statistical regularities effectively act as an implicit language model under greedy decoding. By contrast, Split-and-Share models do not have access to these bilateral dependencies, and ACM further suppresses within-user cues by enforcing lower-order feature reliance. Once an external 6-gram LM is applied, however, its much stronger prior readily compensates for both the lost co-articulatory signal of the Split-and-Share models and the reduced within-user discriminability induced by ACM, thereby revealing the underlying generalization advantage of our approach. Second, all experiments use healthy participants; the extent to which our priors transfer to populations with motor or limb differences (where electrode placement and muscle recruitment differ markedly) remains to be tested.
@@
In sum, we show that the long‑standing vision of “keyboard‑quality” EMG typing can be advanced not only by more data but also by better assumptions. By pairing causal normalization, aggressive yet structured regularization, and a symmetry‑aware encoder, we make significant strides towards truly out‑of‑the‑box wrist‑sEMG typing—paving the way for real‑world assistive and AR/VR interfaces.
7 Acknowledgments
This work was supported by the following awards to JCK: NSF CAREER 1943467, NIH DP2NS122037, NIH R01NS121097. The authors would also like to thank Andrea Ortone for carefully checking the FLOPs calculations and providing helpful feedback.
@@ A.2 Evaluation Details
A.3 Greedy vs. Beam‑Search Decoding
Our main text focuses on beam‑search decoding with an external 6‑gram language model (LM), because this is the configuration most relevant to realistic, latency‑constrained deployment. For completeness, we also report greedy CTC decoding results,i.e. decoding without any LM.
Table 3: Zero-shot CER (%, mean across participants). Columns in gray correspond to training domain validation results, which are reported for transparency but not used as indicators of generalization. Note that the training domain validation results shown here with greedy decoding were computed over all 96 training users; the beam search decoding results are computed on the 18 user subset.
Method Train domain val (greedy) Train domain val (beam)
Other
domain
val
Other
domain
val
Test
domain
val
Test
domain
val
Test
domain
test
Test
domain
test
Sivakumar et al. 2024 22.51 12.14 72.44 72.07 55.57 52.10 55.38 51.78
+ RSG 23.59 13.52 68.55 67.48 52.24 47.26 52.27 47.18
+ RTN 23.49 13.09 64.19 61.95 46.31 39.49 46.07 39.15
+ ACM 34.77 23.47 65.60 63.08 48.87 42.62 49.23 42.62
+ RTN + ACM 32.85 21.71 61.82 58.85 43.72 36.41 43.80 36.42
+ Split 35.59 23.93 61.74 58.64 45.73 37.28 45.69 37.37
+ Share (SplashNet-mini) 37.72 26.44 61.07 58.20 45.33 36.46 45.26 36.41
   + Upscale (SplashNet) 33.57 20.59 60.16 56.95 44.79 35.49 44.78 35.67
Zero‑shot generalization.
Table 3 shows that, in the zero‑shot regime (i.e. on non-training domains), the joint‑hand architecture with RTN and ACM often surpasses the Split‑and‑Share architecture under greedy decoding, even though the ranking reverses once beam search is applied. The most plausible explanation is that the joint‑hand encoder can observe both wrists simultaneously and therefore learns a stronger implicit LM, capturing bi‑ and tri‑gram dependencies spread across hands. While that emergent linguistic prior is helpful when no external LM is available, it becomes redundant—and potentially counter‑productive—once a more reliable 6‑gram LM is introduced at inference time.
Table 4: Finetuned CER (%) with and without beam search.
Method
Test
domain
val
Test
domain
val
Test
domain
test
Test
domain
test
@@
Fine‑tuned models.
A different pattern emerges after user‑specific fine‑tuning (Table 4). None of our variants fully match the baseline of Sivakumar et al. [2024] under greedy decoding, despite decisive gains once beam search is enabled. Because the only changes between the baseline and the "+ RSG" model are (i) four‑fold longer clips during training, (ii) coarser spectral resolution, and (iii) a five‑fold higher probability of channel masking, the degradation must stem from one (or a combination) of these factors.
Two clues implicate aggressive channel masking (ACM). First, the model with RTN but not ACM achieves the best greedy scores, whereas the model with just ACM but not RTN yields the worst greedy performance. Second, the performance gap between the model with just RTN and the model with both RTN and ACM vanishes under beam search, indicating that the external LM compensates for information lost when ACM forces the network to rely on low‑order feature combinations. We therefore hypothesize that ACM, while beneficial for cross‑user generalization, removes within‑user cues that help distinguish confusable keystrokes, a weakness that beam‑search decoding can largely recover.
To confirm that the implicitly increased channel masking in the "+ RSG" model (and all other models) is largely responsible for the uniformly worse greedy decoding performance we see compared to the baseline of Sivakumar et al. [2024], we ran an additional experiment in which we trained a model similar to our +RSG model, but apply SpecAugment on the 33-bin spectrograms from each channel before the bin aggregation of RSG (rather than after). This keeps the extent of masking equivalent to that of Sivakumar et al. 2024 while also allowing us to use the RSG frontend. This model does not perform significantly worse than the baseline of Sivakumar et al. 2024 with greedy decoding, confirming our suspicion that the core reason for the worse greedy decoding results in the finetuned case is increased channel masking.
Table 5: Finetuned CER (%) under greedy decoding. Values are mean $\pm$ standard deviation.
Method
Test
domain
val
Test
domain
test
Sivakumar 2024 $1.39\pm 4.28 11$$1.28\pm 4.45$
+ RSG $12.11\pm 4.67$ $12.50\pm 4.97$
+ RSG w/ pre-aggregation masking $1.02\pm 4.32 11$$1.53\pm 4.70$
In summary, greedy decoding accentuates two complementary inductive biases: (1) joint‑hand encoders learn a useful internal LM, an advantage that vanishes once an external LM is applied, and (2) aggressive masking of input channels trades cross‑user generalization for within‑user discriminability, a trade‑off that RTN and beam‑search decoding can effectively offset.
@@ A.4 Calculation of FLOPs
A.5 Additional Ablations on ACM Masking, RTN Sliding Windows, and RSG
Table 6: Zero-shot CER (%, mean $\pm$ s.d. across participants) for ablations on ACM mask width and RTN sliding window. “-RSG” corresponds to applying ACM on the full-resolution spectrogram as described in the text.
Method
Other
domain
val
Test
domain
val
Test
domain
test
SplashNet-Mini $58.20\pm 10.50$ $36.46\pm 7.09$ $36.41\pm 7.30$
+ 4s SW inference $60.43\pm 7.86$ $37.57\pm 7.18$ $37.71\pm 7.47$
+ 4s SW train + inference $60.43\pm 7.70$ $36.95\pm 7.89$ $36.69\pm 7.80$
+ 16s SW inference $57.24\pm 11.00$ $36.43\pm 7.30$ $36.73\pm 7.40$
+ mask width = 8 $58.12\pm 11.72$ $36.57\pm 7.02$ $36.05\pm 6.74$
+ mask width = 16 $57.65\pm 10.47$ $36.66\pm 7.21$ $36.92\pm 7.40$
- RSG $60.70\pm 11.23$ $37.90\pm 6.58$ $37.19\pm 6.33$
Joint-Hand baseline (+ RSG + RTN + ACM) $58.85\pm 10.50$ $36.41\pm 7.21$ $36.42\pm 7.11$
+ mask width = 8 $59.41\pm 10.63$ $37.55\pm 6.97$ $37.30\pm 6.90$
+ mask width = 16 $57.17\pm 11.32$ $36.95\pm 7.62$ $36.47\pm 7.96$
Table 7: Finetuned CER (%, mean $\pm$ s.d. across participants) for ablations on ACM mask width and RTN sliding window.
Method
Test
domain
val
Test
domain
test
SplashNet-Mini $6.13\pm 2.96$ $5.87\pm 3.04$
+ 4s SW inference $6.10\pm 2.91$ $6.28\pm 3.09$
+ 4s SW train + inference $6.34\pm 2.86$ $6.37\pm 3.10$
+ 16s SW inference $5.76\pm 2.64$ $5.74\pm 2.86$
+ mask width = 8 $5.85\pm 2.72$ $6.02\pm 3.11$
+ mask width = 16 $6.09\pm 2.88$ $5.72\pm 2.96$
- RSG $5.87\pm 2.73$ $5.75\pm 3.06$
Joint-Hand baseline (+ RSG + RTN + ACM) $6.55\pm 2.91$ $6.53\pm 3.27$
+ mask width = 8 $6.14\pm 2.85$ $6.56\pm 3.60$
+ mask width = 16 $6.55\pm 3.09$ $6.45\pm 3.18$
We conducted additional ablations to examine the sensitivity of model performance to the strength of ACM, the presence of RSG, and the temporal window used for RTN normalization. Unless otherwise noted, none of the differences reported here reached significance under a two-tailed paired $t$-test across participants.
ACM mask width.
In our main experiments, ACM uses a maximum frequency mask width of 12 bins, while models without ACM use a maximum width of 4. Here, we trained SplashNet-mini and the Joint-Hand baseline with maximum mask widths of 8 and 16. As shown in Tables 6 and 7, these variations produced small, nonsignificant changes in CER in both the zero-shot and finetuned settings, indicating that performance is not particularly sensitive to the precise mask width within this range.
-RSG.
We also evaluated a variant without RSG, in which ACM was applied to the full-resolution spectrogram by masking a 6-bin dummy vector and inverting the 33-to-6 bin mapping used in RSG to obtain a 33-bin mask. Although performance differences relative to the standard RSG configuration were small and nonsignificant, this variant considerably increases both compute and memory requirements since it has more than sixfold greater input feature dimensionality.
RTN sliding window.
Our original RTN normalization uses all past time points within a sample to compute normalization statistics. We additionally evaluated inference-time sliding-window variants with 4-second and 16-second windows. Using a 4-second window led to a small but statistically significant degradation in zero-shot performance ($p<0.05$, two-tailed), while a 16-second window yielded performance comparable to the default setting. This likely reflects the fact that models were trained on 16-second samples and benefited from matching normalization context at test time. When models were also trained with a 4-second RTN sliding window, this degradation disappeared, and differences were no longer significant in either the zero-shot or finetuned settings.
Summary.
Across all tested configurations, performance differences were small and generally nonsignificant. ACM is robust to maximum frequency mask width changes in the range of 8–16 bins. Removing RSG does not significantly degrade performance but incurs higher compute and memory costs. RTN performance remains stable across temporal windows when training and inference are matched, with only a modest increase in CER when using 4 second windows at inference time alone.
A.6 Aggressive channel masking with mean imputation
Table 8: CER (%, mean $\pm$ s.d. across participants) for different ACM masking configurations.
Method
Test
domain
val
Test
domain
test
Sivakumar et al. 2024 $52.10\pm 5.54$ $51.78\pm 4.61$
+ RSG $47.26\pm 5.26$ $47.18\pm 5.19$
+ ACM $42.62\pm 7.18$ $42.62\pm 7.10$
+ ACM (mean impute) $47.59\pm 7.16$ $48.43\pm 6.24$
+ RTN + ACM $36.41\pm 7.21$ $36.42\pm 7.11$
We performed an ablation to test whether RTN might interact with ACM by stabilizing input features such that the default masking value of 0 corresponds to their per-session mean. When using ACM with RTN, the default mask value of 0 corresponds to the per-sample mean for each feature, whereas under BatchNorm this same value can be far out of distribution for a given feature from a given sample. To test whether it is important that the masking value is the per-sample feature mean, we replaced RTN with standard BatchNorm while setting the ACM masking value to the per-sample mean of each feature, training a Joint-Hand model with RSG, BatchNorm, and ACM using per-channel sample mean imputation.
Somewhat surprisingly, this variant performed on par with the model with RSG alone and substantially worse than the RSG + BatchNorm + ACM configuration with standard zero imputation. This suggests that simply matching the mask value to the per-session mean is not sufficient to reproduce the performance benefits obtained with RSG + RTN + ACM, let alone to obtain the benefits of ACM in the absence of RTN. More generally, these results indicate that the choice of masking value being stable across samples (rather than corresponding to the per-sample mean) appears to be the more important factor for ACM’s effectiveness in this setting.
A.7 UMAP analyses on early intermediate representations
Figure 5: UMAP visualization of model activations after the first TDSConv block for four models (+RSG, +RSG+ACM, +RSG+ACM+RTN, and SplashNet-mini). We extracted activations from every 100th timestep from one session of each of the 8 held-out users. Colors indicate the user identity of each point. In the models without RTN, some users’ representations occupy largely disjoint regions of the activation manifold, whereas models with RTN (+RSG+ACM+RTN and SplashNet-mini) produce markedly more overlapping per-user representations, indicating improved cross-user alignment in the learned feature space.
A.8 Alternative train-test split analyses
Table 9: CER (%, mean $\pm$ s.d. across participants) with a new 78/18 train/validation split. The first column shows performance on the same 18 users when they were included in the original training set (i.e., the previous training domain validation performance), while the remaining columns show results when these 18 users are held out for validation. The test set remains the same 8 held-out users.
Method
Train domain
val (prev. split)
Train domain
val (18 held-out)
Test domain
val (18 held-out)
Test domain
test (18 held-out)
+ RSG $13.52\pm 7.59$ $58.42\pm 10.43$ $47.89\pm 6.36$ $49.01\pm 5.35$
+ RTN $13.09\pm 6.32$ $53.99\pm 11.14$ $42.81\pm 6.75$ $42.62\pm 6.89$
+ ACM $23.47\pm 9.74$ $55.95\pm 10.46$ $44.59\pm 6.47$ $45.10\pm 7.26$
+ RTN + ACM $21.71\pm 9.67$ $51.25\pm 11.49$ $38.72\pm 6.63$ $39.05\pm 7.22$
+ Split $23.93\pm 10.99$ $51.57\pm 11.78$ $39.81\pm 6.72$ $40.14\pm 7.16$
    + Share $26.44\pm 10.64$ $49.84\pm 12.57$ $38.60\pm 7.52$ $38.45\pm 7.86$
A.9 Analyses on emg2pose dataset
Table 10: Mean angular error (degrees) and landmark distance ($m$) under different generalization regimes. All results are in the "tracking" setting.
Generalization Model
Angular
Error
Landmark
Dist
User RTN+ACM 7.5951 10.1567
Stage RTN+ACM 11.1243 15.2404
User, Stage RTN+ACM 10.8433 15.3644
User RTN 7.6007 10.1641
Stage RTN 11.1627 15.2241
User, Stage RTN 10.9085 15.3851
User ACM 7.6388 10.2394
Stage ACM 11.2601 15.4489
User, Stage ACM 11.0343 15.6413
User Baseline 7.6549 10.2585
Stage Baseline 11.2892 15.4360
User, Stage Baseline 11.1222 15.7388
Although the focus of this work is on EMG keystroke decoding, we also performed preliminary experiments to assess how RTN and ACM transfer to the emg2pose benchmark [Salter et al., 2024]. A key consideration is that the baseline model of Salter et al. [2024] uses a learned convolutional featurizer that immediately mixes signals across all electrodes, rather than spectrogram-based per-electrode features as in our keystroke decoding setting. This architectural difference makes a direct application of RTN and ACM less straightforward.
To adapt these methods, we applied RTN and ACM to the outputs of the first convolutional layer in the featurizer. Since these intermediate features are not spectrograms, ACM was implemented by randomly zeroing out 50% of the features in one-third of training samples and 75% in another third. The resulting performance, summarized in Table 10, shows minimal differences relative to the baseline in both mean angular error and landmark distance across user, stage, and combined generalization regimes.
These findings are somewhat unsurprising: applying ACM after early feature mixing likely limits its ability to enforce robust low-order structure, and RTN is less meaningful without clear per-electrode feature boundaries. Future work should explore applying these methods directly to the raw EMG signal before featurization, or using architectures with explicit per-electrode feature streams (e.g., spectrogram-based or structured learned featurizers), where RTN and ACM may have stronger effects.
A.10 Compute Resources